### v2

@@ 7 Conclusion
In this paper we have developed a general framework for mean-field BSDEs whose generators have quadratic growth in the control and its law. Working under the separately quadratic structure (Hq) with bounded terminal condition, we proved global existence and uniqueness of solutions in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, without imposing any smallness condition. The key ingredients were a Malliavin calculus argument for Lipschitz mean-field BSDEs, which yields uniform bounds on the control process, together with BMO estimates and a global stability result that allows us to pass from Lipschitz to quadratic generators. We then extended the analysis to the fully coupled quadratic regime (Hq), in which the generator admit genuine cross-interaction terms. In this more general setting we established existence and uniqueness under a natural smallness condition on the centred terminal condition. As an application, we showed how the fully coupled theory can be used to formulate and solve a mean-field exponential utility maximization problem, thereby generalizing the classical framework of hu2005utility.

### v3

@@ 7 Conclusion
In this paper, we study mean-field BSDEs whose generators of quadratic growth in the control component and in the mean-field term. Our main well-posedness results cover two complementary structural regimes. First, under the unified quadratic condition $(\mathrm{H}_{q}^{\mathrm{unif}})$, we establish global existence and uniqueness of solutions in $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ for bounded terminal conditions of the form $\xi=\phi(W_{t_{1}},\ldots,W_{t_{n}})$, where $\phi$ is bounded and Hölder continuous. Second, under the separably quadratic condition $(\mathrm{H}_{q}^{\mathrm{sep}})$, we prove existence and uniqueness in the same solution class for bounded terminal conditions generated by uniformly continuous functionals of the Brownian path, i.e. $\xi=\phi(W_{[0,T]})$ with $\phi:C([0,T];\mathbb{R}^{d})\to\mathbb{R}$ uniformly continuous in the supremum norm.
Finally, motivated by a mean-field exponential utility maximization problem with market-adaptive constraints, we require a fully coupled quadratic condition allowing cross-interaction terms. For this purpose we work under the quadratic condition $(\mathrm{H}_{q}^{\mathrm{LK}})$ (equivalent to $(\mathrm{H}_{q}^{\mathrm{unif}})$) and develop a dedicated concatenation approach. This yields existence and uniqueness for the associated BSDE under a smallness assumption on the centred terminal condition $\eta:=\xi-\mathbb{E}[\xi]$. As an application, we show that the resulting BSDE characterizes the value function and an optimal strategy, thereby extending the classical framework of Hu et al. (2005) to a mean-field market with constraints driven by an aggregated trading signal.